Operations Report – December 6th

SOL: 9

Name of person filing report: Ashley Kowalski

Non-nominal systems: N/A

Notes on non-nominal systems: N/A


Spirit rover used: Yes
Hours: 203.2 (Start), 203.4 (End)
Beginning charge: 100%
Ending charge: 90%
Currently charging: No

Opportunity rover used: Yes
Hours: 107.3 (Start), 107.9 (End)
Beginning charge: 100%
Ending charge: 69%
Currently charging: No

Curiosity rover used: Yes
Hours: 215.6 (Start), 216.2 (End)
Beginning charge: 100%
Ending charge: 63%
Currently charging: No

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: ~350 gallons used to-date
Water (static tank): ~200 gallons (This is my best guess, but I am requesting Mission Support to let us know a more accurate value, if possible.)
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer replaced the battery in EVA Suit #6 today. She wore this suit out on the long afternoon EVA today, and the suit’s air flow system worked well. Please let us know what to do with the old battery.

Additionally, the Crew Engineer’s headset (headset #4) seemed to be working normally today and had no issues. However, the GreenHab Officer’s headset (headset #3) seemed to be experiencing issues today. The GreenHab Officer could hear his crewmates speaking, but his crewmates could not hear him. We switched his headset and he used headset #5 for his EVA today instead. We will monitor and report back on its status.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~44 gallons (~5 gallons today)
Heater: On
Supplemental light: On
Harvest: N/A

Summary of ScienceDome operations: N/A
Dual split: Per Mission Support’s message, the heat is on in the ScienceDome and set to 65 F.

Summary of RAM operations: No RAM operations today.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: The Crew Engineer replaced the heat furnace air filter today. Also, the Crew Engineer acknowledges the email that was sent to the crew yesterday regarding batteries for the walkie talkies. The large sharpies work well. Finally, the toilet seems to now be functioning well. No problems today.

Copyright © The Mars Society. All rights reserved. | Main Site